# Decision-making with Speculative Opponent Models

### 06 Mar 2024 v2

@@ Appendix B Concerns about POMDP setting
Appendix C Environmental settings
Predator-prey: The states, observations, actions, state transition function, and reward function of each agent is formulated below by following the POMG convention.
  • The states and observations. A grid world of size $x\times x$, e.g. Figure 2(a) (in the main paper) is a state of size $7\times 7$ containing four predators and two preys. The observation of agent $i$ is the coordinates of its location, its ID, and the coordinates of the prey $k$ relative to $i$ in $l\times l$ view, if observed.
@@ Appendix C Environmental settings
  • State transition $\mathcal{T}$. The new state after the transition is the map with updated positions of all agents due to agents moving in the grid world. The termination condition for this task is when all preys are dead or for 100 steps.
  • Rewards $\mathcal{R}$. All agents move within the map and can not exceed the boundary. Since the predators cooperate with each other, they share the team reward. The predators share a reward of 5 if two or more of them catch the prey simultaneously, while they are given a negative reward of -0.5 if only one predator catches the prey.
@@
  • The states and observations. At each time step, agents get local observations within their field of view $5\times 5$, which contains information (board, position,ammo) about the map. The agent obtain the information of the Blast Strength, whether the agent can kick or not, the ID of their teammate and enemies, as well as the agent’s current blast strength and bomb life.
@@
Appendix D Evaluation results
@@
The number of quantiles may influence the algorithm performance, for example, using too few quantiles can lead to poor performance. To further reveal the influence of the quantiles number $K$ on learning the centralized distributional critic, we conduct another ablation study for different values of $K$. The results are shown in Figure A2(d), A2(e) and A2(f) where we can see that the performance of $K=5$ quantiles is consistently better than that of $K=3$ quantiles for the entire training period. This is aligned with the intuition that more quantiles can support more fine-grained distribution modeling. Thus, it better captures the return randomness and improves performance. However, the running time would increase as the number of quantiles becomes larger. To balance computational overhead and the performance, we set $K=5$ throughout the experiments. Finally, we also present the additional experiment results of ablation studies in Figure A1 and Figure A2. The results show that DOMAC outperforms baseline methods as well in other scenarios, which is supported by the results that DOMAC has the lowest entropy in Figure A1(a), A1(b),A1(c) and the highest average episode rewards in Figure A2(a), A2(b) and A2(c).
(a) Agent entropy for PP-2v1.
(b) Agent entropy for Pomm-FFA.
(c) Agent entropy in Pomm-Team.
(d) OM accuracy for PP-2v1.
@@
Figure A1: Impacts of the distributional critic on opponent models (OM) and policies for the rest games.
(a) Study of OM in PP-2v1.
(b) Study of OM in Pomm-FFA.
(c) Study of OM in Pomm-Team.
(d) Study for K in PP-2v1.
(e) Study for K in Pomm-FFA.
(f) Study for K in Pomm-Team.
Figure A2: Ablation studies for opponent models (OM) and quantile number $K$ in PP-2v1, Pomm-FFA, Pomm-Team.

### 22 Mar 2024 v3

@@ Appendix A Distributional opponent model aided policy gradient theorem
Appendix B Environmental settings
Predator-prey: The states, observations, actions, and state transition function of each agent is formulated as below.
  • The observations. Observations consist of high-level feature vectors containing relative distances to other agents and landmarks.
@@ Appendix B Environmental settings
  • State transition $\mathcal{T}$. The new state after the transition is the map with updated positions of all agents due to agents moving in the world. The termination condition for this task is when all preys are dead or for 500 steps.
@@
  • The states and observations. At each time step, agents get local observations within their field of view $5\times 5$, which contains information (board, position,ammo) about the map. The agent obtain information of the Blast Strength, whether the agent can kick or not, the ID of their teammate and enemies, as well as the agent’s current blast strength and bomb life.
@@
SMAC Setting:
  • The states and observations. At each time step, agents receive local observations, which contain information (distance, relative x, relative y, health, shield, and unit type) about the map within a circular area for both allied and enemy units and makes the environment partially observable for each agent. The global state is composed of joint observations, which could be used during training. All features, both in the global state and in individual observations of agents, are normalized by their maximum values.
[…]
  • Rewards. At each time step, the agents receive a joint reward equal to the total damage dealt to the enemy agents. In addition, agents receive a bonus of 10 after killing each opponent, and 200 after killing all opponents for winning the game. The rewards are normalized to ensure that the maximum cumulative reward achievable in each scenario is around 20.
Appendix C Evaluation results
Analysis of the computation complexity
@@ Appendix C Evaluation results
(a) Study of OM in PP-3v1.
(b) Study of OM in Pomm-FFA.
(c) Study of OM in Pomm-Team.
(d) OM accuracy for PP-3v1.
@@
(j) Study for K in PP-3v1.
(k) Study for K in Pomm-FFA.
(l) Study for K in Pomm-Team.
Figure A1: Ablation studies for opponent models (OM) and quantile number $K$ in PP-3v1, Pomm-FFA, Pomm-Team.
(a) 3v1
(b) FFA
(c) 2s3z
Figure A2: Average entropy of policies for game FFA and 2s3z.